# Rendering Wireless Environments Useful for Gradient Estimators: A Zero-Order Stochastic Federated Learning Method

### v2

@@ C-B Proof of Theorem 1: Convergence Analysis
From Lemma 2 and by looking closely at the use of the local functions’ boundedness property in (9), we can say $\|g_{k}\|^{2}\leq c$ almost surely for some $c>0$. Thus, again by Assumption 5, $\lim_{K\rightarrow\infty}\sum_{k=0}^{K}\alpha_{k}^{2}<\infty.$ We conclude that
@@
$\begin{split}\|\nabla F(\theta_{k_{l}+1})\|&\geq\|\nabla F(\theta_{k_{l}})\|-\|\nabla F(\theta_{k_{l}+1})-\nabla F(\theta_{k_{l}})\|\\ &\geq\rho-\varepsilon-L\|\theta_{k_{l}+1}-\theta_{k_{l}}\|\\ &=\rho-\varepsilon-L\alpha_{k_{l}}\|g_{k_{l}}\|\\ &\geq\rho-\varepsilon-L\sqrt{c}\alpha_{k_{l}},\end{split}$
Since $k_{l}\rightarrow\infty$ as $l\rightarrow\infty$, we can always find a subsequence of $(k_{l_{p}})_{p\in\mathbb{N}}$ such that $k_{l_{p+1}}-k_{l_{p}}>1$. As $\alpha_{k_{l}}$ is vanishing, we consider $(k_{l})_{l\in\mathbb{N}}$ starting from $\alpha_{k_{l}}<\frac{\rho-\varepsilon}{L\sqrt{c}}$. Thus,
$\begin{split}\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\|\nabla F(\theta_{k+1})\|^{2}&\geq(\rho-\varepsilon)^{2}\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\\ &-2(\rho-\varepsilon)L\sqrt{c}\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\alpha_{k}\\ &+L^{2}c\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\alpha_{k}^{2}\\ &>+\infty.\end{split}$
On the R.H.S., the first series diverges, and the second and the third converge by Assumption 5. This implies that $\sum_{k}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}$ diverges. This is a contradiction as this series converges almost surely by (15). Therefore, hypothesis (H) cannot be true and $\|\nabla F(\theta_{k})\|$ converges to zero almost surely.

### v3

@@ C-B Proof of Theorem 1: Convergence Analysis
To prove the finiteness of $\sum_{k=0}^{K}\alpha_{k}^{2}\|g_{k}\|^{2}$, we let $X_{k}$ be a centered Gaussian process. We know that for any $\sigma^{2}$-subgaussian random variables $X_{1},\ldots,X_{K}$, we have
$\mathbb{P}\left[\sup_{1\leq k\leq K}X_{k}\geq\sqrt{2\sigma^{2}(\log K+t)}\right]\leq e^{-t}$
since let $u:=\sqrt{2\sigma^{2}(\log K+t)}$,
$\begin{split}\mathbb{P}\left[\sup_{1\leq k\leq K}X_{k}\geq u\right]&=\mathbb{P}\left[\exists k,X_{k}\geq u\right]\\ &\leq\sum_{k=1}^{K}\mathbb{P}\left[X_{k}\geq u\right]\leq Ke^{-\frac{u^{2}}{2\sigma^{2}}}=e^{-t}.\end{split}$
Then, for $t=c\log K$ with $c>1$, $\mathbb{P}\left[\sup_{1\leq k\leq K}X_{k}\geq\sqrt{2\sigma^{2}(1+c)\log K}\right]\leq\frac{1}{K^{c}}$ and
$\sum_{K=1}^{\infty}\mathbb{P}\left[\sup_{1\leq k\leq K}X_{k}\geq\sqrt{2\sigma^{2}(1+c)\log K}\right]\leq\sum_{K=1}^{\infty}\frac{1}{K^{c}}<\infty.$
By Borel-Cantelli Lemma, we have
$\mathbb{P}\left(\lim_{K\rightarrow\infty}\sup\{\sup_{1\leq k\leq K}X_{k}\geq\sqrt{2\sigma^{2}(1+c)\log K}\}\right)=0.$
Then, w.p.1 $\exists K^{\prime}<\infty$ s.t. $\forall K\geq K^{\prime}$, $\sup_{1\leq k\leq K}X_{k}\leq\sqrt{2\sigma^{2}(1+c)\log K}$ with $c>1$.
Thus, the supremum of the Gaussian elements of $\|g_{k}\|$, i.e., $h_{i,k+1},h_{i,k}$, $n_{i,k}$, and $n_{i,k+1}$, at worst grow as $O(\sqrt{log(k+1)})$, and hence the upper bound on $\|g_{k}\|$ grows as $c_{4}\sqrt{\log(k+1)}$, where $c_{4}=\beta_{2}N\big{(}\frac{C}{\sigma_{h}}+\sigma_{n}\big{)}\sqrt{2(1+c)}$:
$\begin{split}\|g_{k}\|&=\Big{\|}\Phi_{k}\Big{(}\sum_{i=1}^{N}h_{i,k+1}\frac{f_{i}\big{(}\theta_{k}^{\prime},\xi_{k+1}\big{)}}{\sigma_{h}^{2}}+n_{i,k+1}\Big{)}\Big{\|}\\ &\leq\|\Phi_{k}\|\Big{(}\sum_{i=1}^{N}\Big{\|}h_{i,k+1}\frac{C}{\sigma_{h}^{2}}\Big{\|}+\|n_{i,k+1}\|\Big{)}\\ &\leq\beta_{2}N\bigg{(}\frac{C}{\sigma_{h}^{2}}\sqrt{2\sigma_{h}^{2}(1+c)\log(k+1)}\\ &\hskip 56.9055pt+\sqrt{2\sigma_{n}^{2}(1+c)\log(k+1)}\bigg{)}\\ &=\beta_{2}N\big{(}\frac{C}{\sigma_{h}}+\sigma_{n}\big{)}\sqrt{2(1+c)\log(k+1)}\end{split}$
Then, we write $\sum_{k=0}^{K}\alpha_{k}^{2}\|g_{k}\|^{2}=\sum_{k=0}^{K^{\prime}}\alpha_{k}^{2}\|g_{k}\|^{2}+\sum_{k=K^{\prime}}^{K}\alpha_{k}^{2}\|g_{k}\|^{2}$, where $\sum_{k=0}^{K^{\prime}}\alpha_{k}^{2}\|g_{k}\|^{2}<\infty$ for $K^{\prime}<\infty$.
We know that $\forall\epsilon>0$, $\log(k+1)\leq\frac{1}{\epsilon}(k+1)^{\epsilon}$. Thus, by Assumption 5, for $2\upsilon_{1}=1+\epsilon^{\prime}$,
$\begin{split}&\lim_{K\rightarrow\infty}\sum_{k=K^{\prime}}^{K}\alpha_{k}^{2}\|g_{k}\|^{2}\\ &\leq\lim_{K\rightarrow\infty}c_{4}^{2}\sum_{k=K^{\prime}}^{K}\alpha_{k}^{2}\log(k+1)\\ &\leq\lim_{K\rightarrow\infty}c_{4}^{2}\sum_{k=K^{\prime}}^{K}\frac{1}{(k+1)^{1+\epsilon^{\prime}}}\times\frac{1}{\epsilon}(k+1)^{\epsilon}<\infty,\\ &\hskip 142.26378pt\;\;\forall\epsilon^{\prime}>\epsilon>0.\end{split}$ (15)
We conclude that
@@
$\begin{split}\|\nabla F(\theta_{k_{l}+1})\|&\geq\|\nabla F(\theta_{k_{l}})\|-\|\nabla F(\theta_{k_{l}+1})-\nabla F(\theta_{k_{l}})\|\\ &\geq\rho-\varepsilon-L\|\theta_{k_{l}+1}-\theta_{k_{l}}\|\\ &=\rho-\varepsilon-L\alpha_{k_{l}}\|g_{k_{l}}\|\\ &\geq\rho-\varepsilon-Lc_{4}\sqrt{\log(k_{l}+1)}\alpha_{k_{l}},\end{split}$ (17)
Since $k_{l}\rightarrow\infty$ as $l\rightarrow\infty$, we can always find a subsequence of $(k_{l_{p}})_{p\in\mathbb{N}}$ such that $k_{l_{p+1}}-k_{l_{p}}>1$. As $\alpha_{k_{l}}$ is vanishing, we consider $(k_{l})_{l\in\mathbb{N}}$ starting from $\alpha_{k_{l}}<\frac{\rho-\varepsilon}{Lc_{4}\sqrt{\log(k_{l}+1)}}$. As a reminder, we can always find $\epsilon>0$ such that $\log(x)\leq\frac{1}{\epsilon}x^{\epsilon}$. Thus,
$\begin{split}&\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\|\nabla F(\theta_{k+1})\|^{2}\\ \geq&(\rho-\varepsilon)^{2}\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\\ &-2(\rho-\varepsilon)Lc_{4}\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\alpha_{k}\sqrt{\log(k+1)}\\ &+L^{2}c_{4}^{2}\sum_{k=0}^{\infty}\alpha_{k+1}\gamma_{k+1}\alpha_{k}^{2}\log(k+1)\\ >&+\infty.\end{split}$ (18)
as the first series diverges, and the second and the third converge by Assumption 5. This implies that the series $\sum_{k}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}$ diverges. This is a contradiction as this series converges almost surely by (16). Therefore, hypothesis (H) cannot be true and $\|\nabla F(\theta_{k})\|$ converges to zero almost surely.